OpenStudy (annipuppi):

Complete the square 1. x^2+9x+ 2. x^2+22x+ 3. x^2+30x+

OpenStudy (anonymous):

what is the last term of the question?

OpenStudy (annipuppi):

I don't know... that's all it gives me

OpenStudy (annipuppi):

#3 should be x^2-30x instead though

OpenStudy (anonymous):

ok lets do that first

OpenStudy (anonymous):

takes only two steps 1) what is half of 30?

OpenStudy (annipuppi):

15

OpenStudy (anonymous):

ok good part 2) what is \(15^2\) ?

OpenStudy (anonymous):

225

OpenStudy (annipuppi):

225

OpenStudy (anonymous):

then that is your answer \(225\) and \[x^2-30x+225=(x-15)^2\]

OpenStudy (annipuppi):

so is a 20.25 and b 121?

OpenStudy (anonymous):

b is 121 yes

OpenStudy (anonymous):

and a is \(\frac{81}{4}\) which is probably what you wrote as a decimal

OpenStudy (annipuppi):

okay thank you! That helped a lot

OpenStudy (anonymous):

yw, you have to say it is easy once you know what to do right? some people treat this completing the square as if it were rocket science, but it only takes two steps